OpenStudy (anonymous):

What is the solution to the proportion? 8/x=4/5 A:2.5 B:5 C:9 D:10

OpenStudy (anonymous):

4/5 is equal to 8/10. So find x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\frac{ 4 }{ 5 }\times \frac{ 2 }{ 2 }=\frac{ 8 }{ 10 }\]

OpenStudy (anonymous):

thank u @abdul_shabeer and @RoseDryer